OpenStudy (anonymous):

help me please!! what is the product of the radical expression? (7- sqr root 2)(8+ sqr root 2) a. 13+15 sqr root 2 b. 54- sqr root 2 d. 54+56 sqr root 2

OpenStudy (yttrium):

You know something about foil method?

OpenStudy (anonymous):

no

OpenStudy (nirmalnema):

\[(7-\sqrt{2})(8+\sqrt{2})\]

OpenStudy (tkhunny):

Perfect. Never learn it. Just learn to multiply and remember the distributive property. \((7- \sqrt{2})(8+ \sqrt{2})\) Using the distributive property, we have: \(7\cdot(8+ \sqrt{2}) - \sqrt{2}\cdot(8+ \sqrt{2})\) Using the distributive property, we have: \(7\cdot 8 + 7\cdot\sqrt{2} - \sqrt{2}\cdot 8 - \sqrt{2}\cdot\sqrt{2})\) Multiply all the pieces and see where you can simplify it!

OpenStudy (nirmalnema):

\[56+7\sqrt{2}-8\sqrt{2}-2\]

OpenStudy (nirmalnema):

\[54+\sqrt{2}(7-8)\]

OpenStudy (nirmalnema):

\[54-\sqrt{2}\]

OpenStudy (yttrium):

We're just applying same approach. Anway it will lead to 54 - sqrt(2)

OpenStudy (anonymous):

thank you can you help me with some other problems as well?

OpenStudy (nirmalnema):

yup sure..i will try

OpenStudy (yttrium):

Of course,.:)

OpenStudy (anonymous):

what is the simplest form of the expression? sqr root 6x^8y^9/ sqr root 5x^2y^4

OpenStudy (yttrium):

But you better make it as anew post. :))